Square root of 6 * square root of 5 * square root of 4 * square root of 3 * square root of 2 * square root of 1 = k square root of 5
What is the value of k?
Thnx

Answer by rothauserc(4718)

Square root of 6 * square root of 5 * square root of 4 * square root of 3 * square root of 2 * square root of 1 = k square root of 5
we need the prime factors for each term
(3^1/2 * 2^1/2) * 5^1/2 * (2^1/2 * 2^1/2) * 3^1/2 * 2^1/2 * 1^1/2
combine like terms by adding their exponents and note 1^1/2 = 1
3 * 2^2 * 5^1/2 = 12 * 5^1/2
k = 12